7. How many mL of a 0.90 % (m/v) solution of NaCl are needed to provide 45 mg of NaCl?
Answer #1

Given:
(m/v)% = 0.90 %
Mass of NaCl = 45 mg = 0.045 g

Now use:
(m/v)% = mass of NaCl in g * 100 / volume in mL
0.90 = 0.045 * 100 / volume in mL
volume = 5.0 mL

Answer: 5.0 mL